# LW-FedSSL: Resource-Efficient Layer-wise Federated Self-Supervised Learning

### v2

@@ Abstract
Many studies integrate federated learning (FL) with self-supervised learning (SSL) to take advantage of raw training data distributed across edge devices. However, edge devices often struggle with high computation and communication costs imposed by SSL and FL algorithms. To tackle this hindrance, we propose LW-FedSSL, a layer-wise federated self-supervised learning approach that allows edge devices to incrementally train a single layer of the model at a time. Our LW-FedSSL comprises server-side calibration and representation alignment mechanisms to maintain comparable performance with end-to-end federated self-supervised learning (FedSSL) while significantly lowering clients’ resource requirements. In a pure layer-wise training scheme, training one layer at a time may limit effective interaction between different layers of the model. The server-side calibration mechanism takes advantage of the resource-rich server in an FL environment to ensure smooth collaboration between different layers of the global model. During the local training process, the representation alignment mechanism encourages closeness between representations of FL local models and those of the global model, thereby preserving the layer cohesion established by server-side calibration. Our experiments show that LW-FedSSL has a $\times3$ lower memory requirement and a 3.2×3.2$\times$3.2 × cheaper communication cost than its end-to-end counterpart. We also explore a progressive training strategy called Prog-FedSSL that outperforms end-to-end training with a similar memory requirement and a 1.8×1.8$\times$1.8 × cheaper communication cost.
Index Terms:
@@
I Introduction
@@
A significant portion of real-world data, valuable for practical AI applications, resides on edge devices. Such data is often unlabeled, making the adoption of self-supervised learning (SSL) strategies [1, 2, 3] essential. SSL enables models to learn from raw data by generating their own supervisory signals, eliminating the need for labels. Most SSL methods follow centralized training schemes, often requiring an exhaustive data collection process to build a centrally stored dataset. Although some data, such as natural images, can be conveniently collected, it can raise significant concerns for privacy-sensitive data, such as medical data. Meanwhile, decentralized learning approaches like federated learning (FL) [4, 5] enables privacy-preserving collaborative model training. An FL system functions by collecting trained model parameters from edge devices (a.k.a., clients) rather than data. This mechanism safeguards the confidentiality of private data held by different parties. As a result, federated self-supervised learning (FedSSL) approaches have recently emerged as a potential strategy for leveraging raw data in distributed environments [6, 7, 8].
@@
Self-supervised learning (SSL), especially when involving state-of-the-art models like transformers [9, 10], can impose substantial computation demands. Unfortunately, clients in an FL environment often operate with limited computation and communication resources. Therefore, it is crucial to explore resource-efficient SSL approaches for FL systems. A straightforward way to reduce computation demands on low-memory devices is to conduct the training with a smaller batch size. However, using a small batch size for SSL approaches can diminish the quality of learned representations and compromise performance [1, 7]. Moreover, simply reducing the batch size does not lower the FL communication cost since clients still need to exchange the full model with the server. On the other hand, using a large batch size would exclude low-memory devices from participating in the FL process.
One promising solution is the layer-wise training approach, where a model is systematically trained one layer at a time, gradually increasing its depth. The term “layer” can refer to either an individual layer or a block comprising multiple layers. In layer-wise training, only a single layer is active at a time for training, significantly reducing computation requirements on FL clients. Moreover, only the active layer needs to be exchanged with the server, further reducing communication costs. Table I shows a significant reduction in a client’s resource requirements for layer-wise training compared to end-to-end training.
Although layer-wise training significantly reduces resource requirements, it may not be able to match the performance of end-to-end training as shown in Table I. In each stage of layer-wise training, a new layer is sequentially attached to the model, and all preceding layers are frozen, excluding them from further updates. Given a fixed number of FL rounds for each stage, it can be challenging to train the newly added layer to work seamlessly with the previous layers, especially if the number of rounds is small. Although increasing the number of rounds could potentially alleviate this problem, it would impose extra communication and computation overhead on the clients, canceling the benefits of layer-wise training.
To address these challenges, we propose LW-FedSSL, by introducing two simple, yet effective mechanisms: server-side calibration and representation alignment. Although clients in an FL environment may have limited resources, the central server typically possesses significant computation power. Therefore, in LW-FedSSL, the server-side calibration mechanism leverages the computation resources of the central server to train the global model in an end-to-end manner using an auxiliary dataset. This mechanism solely operates on the central server and incurs no additional overhead on the clients. Meanwhile, the representation alignment mechanism encourages alignment between local models and the server-side-trained global model during the local training process.
Our contributions can be summarized as follows: (i) We conduct an in-depth exploration of layer-wise training for FedSSL, considering various empirical aspects. (ii) We introduce LW-FedSSL, a novel layer-wise training framework designed to improve the resource efficiency of SSL within the FL paradigm. LW-FedSSL can compete with conventional end-to-end training while significantly lowering the resource requirements of FL clients. (iii) We also explore a progressive training approach called Prog-FedSSL. Although the progressive approach requires more resources than the layer-wise approach, it still remains more resource-efficient than end-to-end training while significantly outperforming both.
@@
TABLE I: FedMoCo (end-to-end) vs. LW-FedMoCo (layer-wise). We use a ViT-Tiny backbone with MoCo v3 as the SSL approach. The FL process involves 10 clients using a batch size of 1024 and the number of communication rounds set to 180. The unlabeled portion of STL-10 dataset is used to setup the clients’ data, while the labeled training and testing portions are used for linear evaluation. We compare resource requirements for a single client.
@@
FedMoCo 15601 7352.49 54.99
LW-FedMoCo 3927 613.19 46.41
II Preliminary
In this section, we present fundamental concepts related to federated learning and self-supervised learning.
II-A Federated Learning
A typical FL [5] process involves a central coordinating server and a set of client devices, each of which stores private data. The goal of FL is to train a global model $M$, by learning from the data residing on clients while ensuring their privacy. The FL objective can be expressed as
@@ II-A Federated Learning
where $N$ is the number of clients and $\mathcal{L}$ is the local loss. Here, $w^{i}=|D^{i}|/|D|$, which is the weight term of $i$-th client, with $|D^{i}|$ denoting the number of samples in local dataset $D^{i}$ and $|D|$ denoting the total number of samples in ${D=\bigcup_{i=1}^{N}D^{i}}$. As shown in Fig. 1, a single FL communication round consists of the following four key steps. At the beginning of each round, (i) the server distributes a base global model $M$ to all participating clients. (ii) Each client $i\in[1,N]$ performs training on its local dataset $D^{i}$ to produce a local model $M^{i}$. After the training, (iii) local models are transmitted back to the server. Finally, (iv) the sever aggregates the received local models through weighted averaging [5] to update the global model: $M\leftarrow\sum^{N}_{i=1}w^{i}M^{i}$. These four steps are repeated for a number of rounds $R$.
Figure 1: Four key steps in an FL communication round.
Figure 2: Self-supervised learning with MoCo v3. We omit the negative samples for clarity.
Figure 3: Local training process of LW-FedMoCo across different stages for $i$-th client. Each stage $s\in[1,S]$ increases the depth of the encoder $F$ by adding a new layer $L_{s}$. During stage $s$, only the corresponding layer $L_{s}$ is actively updated, while prior layers (i.e. $L_{1}$ to $L_{s-1}$) are frozen.
II-B Self-Supervised Learning
Self-supervised learning (SSL) offers a means to utilize unlabeled data, which is often more readily available than labeled data, for model training. SSL generates its own supervisory signals to learn valuable representations from the unlabeled data. We select MoCo v3 [11] as the SSL approach in our work. As shown in Fig 2, MoCo v3 features a Siamese network structure with an actively trained online branch, and a target branch, which is a moving-averaged copy of the online branch. The online branch consists of an encoder $F$, a projection head $H$, and a prediction head $P$. Meanwhile, the target branch consists of a momentum encoder $F_{k}$ and a momentum projection head $H_{k}$. Given an input sample $x$, it undergoes augmentation, creating two views, $x_{q}$ and $x_{k^{+}}$. These views are then fed into the online and target branches, respectively, to obtain a vector $q$ and its corresponding positive pair, $k^{+}$. Likewise, negative pairs, $k^{-}$, are derived from other samples within the same batch through the target branch. The online branch is trained using the InfoNCE loss [12],
@@ II-B Self-Supervised Learning
II-C Federated Self-Supervised Learning
Although numerous studies on FL primarily focus on supervised learning, this means that device owners are tasked with labeling data, which is a tedious process. Consequently, client devices often lack labeled data due to various reasons, including limitations in data collection methods, labeling constraints, and insufficient annotation expertise and resources. In response, federated self-supervised learning (FedSSL) [13, 14, 15] emerges as a solution to extend the applicability of FL to new applications and use cases in domains where labeled data is scarce or difficult to acquire. Therefore, in the context of FedSSL, the local dataset $D^{i}$ within a client $i$ is unlabeled, and thus, the local training process involves utilizing a form of SSL loss to train the local model $M^{i}$ on $D^{i}$. When using MoCo v3 as the SSL approach, this local model can comprise an encoder $F^{i}$ and its accompanying MLP heads, $H^{i}$ and $P^{i}$.
III Related Work
@@
Some studies [20, 21] reduce computation costs of FL clients by selectively dropping parts of the model. Federated Dropout [20] randomly removes a subset of neurons or filters from a layer to reduce the computation cost while relying on compression techniques to minimize the communication overhead. FjORD [21], a similar approach, introduces an ordered dropout mechanism instead of random dropping. Meanwhile, [18] proposes a progressive training approach for FL called ProgFed. While similar to layer-wise training in gradually increasing the model depth, ProgFed differs by training all existing layers at each stage. All aforementioned studies primarily focus on federated supervised learning tasks that require labeled data for training.
On the other hand, [22] studies layer-wise SSL in the speech domain. While the study was intended for an on-device training scenario, the authors mainly examined it in a centralized setting. Another study in [19] also explores layer-wise training for FedSSL, in which the authors propose a depth dropout technique which aims to reduce both computation and communication overhead but may lead to a slight decrease in performance. In contrast to the above studies, our work offers an in-depth exploration of layer-wise training in the context of FedSSL. Moreover, as we have previously demonstrated, layer-wise training alone may not achieve the same level of performance as end-to-end training. To bridge this gap, we introduce LW-FedSSL, a layer-wise training approach that can compete end-to-end training with significantly lower client resource requirements. In addition, we explore a progressive training approach called Prog-FedSSL, which outperforms end-to-end training while requiring fewer resources.
@@
11:         $L_{s}^{i},H^{i},P^{i}\leftarrow\text{Local\_SSL}(i,F,H,P)$ Algo. 2
@@
IV Proposed Method
IV-A LW-FedMoCo
Firstly, we discuss the pure layer-wise training baseline, referred to as LW-FedMoCo. Fig. 3 illustrates the local training process of LW-FedMoCo across different stages. For an encoder $F$ with a total of $S$ layers, the training process is divided into $S$ stages. Each stage $s\in[1,S]$ runs for a fixed number of FL communication rounds, $R_{s}$. The training process starts with an empty encoder $F$ without any layers, and each stage $s$ sequentially adds a new layer $L_{s}$ to $F$, gradually increasing its depth. During stage $s$, only the corresponding layer $L_{s}$ is updated, while all prior layers $L_{1}$ to $L_{s-1}$ within $F$ are frozen, serving only for inference. This strategy saves memory and computation, benefiting resource-constrained devices in FL environments. Moreover, in each communication round, clients are only required to send the active layer $L_{s}$ (and the MLP heads) to the server for aggregation. Likewise, the server only needs to broadcast the aggregated layer $L_{s}$ back to clients. This reduces both upload and download communication costs for clients. Although layer-wise training offers better resource efficiency than end-to-end training, there remains a notable performance gap in linear evaluation, as we have shown in Table I.
@@ IV-A LW-FedMoCo
Input: encoder $F$, projection head $H$, prediction head $P$, local dataset $D^{i}$, number of local epochs $E$, momentum $\mu$, temperature $\tau$, weight term $\alpha$
@@
1:Client executes:
2:Local_SSL($i,F,H,P$):
3:Online branch: $F^{i}\leftarrow F$ and $H^{i}\leftarrow H$ and $P^{i}\leftarrow P$
4:Target branch: $F^{i}_{k}\leftarrow F$ and $H^{i}_{k}\leftarrow H$
5:for epoch $e=1,2,\dots,E$ do
6:     for each batch $x\in D^{i}$ do
7:         $x_{1}\leftarrow\text{{Augment}}(x)$ and $x_{2}\leftarrow\text{{Augment}}(x)$
8:         $z_{1}\leftarrow F(x_{1})$ and $z_{2}\leftarrow F(x_{2})$
9:         $z_{1}^{i}\leftarrow F^{i}(x_{1})$ and $z_{2}^{i}\leftarrow F^{i}(x_{2})$
10:         $q_{1}^{i}\leftarrow P^{i}(H^{i}(z_{1}^{i}))$ and $q_{2}^{i}\leftarrow P^{i}(H^{i}(z_{2}^{i}))$
11:         $k_{1}^{i}\leftarrow H^{i}_{k}(F^{i}_{k}(x_{1}))$ and $k_{2}^{i}\leftarrow H^{i}_{k}(F^{i}_{k}(x_{2}))$
12:         $\ell_{\text{con}}\leftarrow\ell(q^{i}_{1},k^{i}_{2},\tau)+\ell(q^{i}_{2},k^{i}_{1},\tau)$ Eq. (2)
13:         $\ell_{\text{align}}\leftarrow\ell(z^{i}_{1},z_{2},\tau)+\ell(z^{i}_{2},z_{1},\tau)$ Eq. (3)
14:         $\mathcal{L}\leftarrow\ell_{\text{con}}+\alpha\ell_{\text{align}}$
15:         $F^{i},H^{i},P^{i}\leftarrow\text{Update}(\nabla\mathcal{L})$
16:         $F^{i}_{k},H^{i}_{k}\leftarrow\text{Momentum\_Update}(F^{i},H^{i},\mu)$
17:$L_{s}^{i}\leftarrow\text{Get active layer from }F^{i}$
18:return $L_{s}^{i},H^{i},P^{i}$
TABLE II: Given an encoder $F$ with a total of $S$ layers, we compare the characteristics of different approaches at stage $s\in[1,S]$.
@@
LW-FedMoCo layer-wise $S$ $L_{s}$ $L_{1}$ to $L_{s-1}$ $L_{s}$ $L_{s}$ $L_{s}$

### v3

@@ Abstract
Many studies integrate federated learning (FL) with self-supervised learning (SSL) to take advantage of raw data distributed across edge devices. However, edge devices often struggle with high computation and communication costs imposed by SSL and FL algorithms. To tackle this hindrance, we propose LW-FedSSL, a layer-wise federated self-supervised learning approach that allows edge devices to incrementally train a single layer of the model at a time. We introduce server-side calibration and representation alignment mechanisms to ensure LW-FedSSL delivers performance on par with conventional federated self-supervised learning (FedSSL) while significantly lowering resource demands. In a pure layer-wise training scheme, training one layer at a time may limit effective interaction between different layers of the model. The server-side calibration mechanism takes advantage of the resource-rich FL server to ensure smooth collaboration between different layers of the global model. During local training, the representation alignment mechanism encourages closeness between representations of local models and those of the global model, thereby preserving the layer cohesion established by server-side calibration. With the proposed mechanisms, LW-FedSSL achieves a $\times3$ reduction in memory usage, 2.1×2.1$\times$2.1 × fewer computational operations (FLOPs), and a 3.2×3.2$\times$3.2 × lower communication cost while maintaining the same level of performance as its end-to-end training counterpart. Additionally, we explore a progressive training strategy called Prog-FedSSL, which matches end-to-end training in memory requirements but offers a 1.8×1.8$\times$1.8 × reduction in FLOPs and communication costs. Although Prog-FedSSL is not as resource-efficient as LW-FedSSL, its performance improvements make it a suitable candidate for FL environments with more lenient resource constraints.
keywords:
@@
journal: Information Fusion
[…]
[label1]organization=Department of Computer Science and Engineering, Kyung Hee University, city=Yongin-si, state=Gyeonggi-do 17104, country=South Korea
[…]
[label2]organization=Vietnam - Korea University of Information and Communication Technology, city=Danang, country=Vietnam
1 Introduction
@@
A significant portion of real-world data, valuable for practical AI applications, resides on edge devices. Such data is often unlabeled, making the adoption of self-supervised learning (SSL) strategies [1, 2, 3] essential. SSL enables models to learn from raw data by generating their own supervisory signals, eliminating the need for labels. Most SSL methods follow centralized training schemes, often requiring an exhaustive data collection process to build a centrally stored dataset. Although some data, such as natural images, can be conveniently collected, it can raise significant concerns for privacy-sensitive data, such as medical data. Meanwhile, decentralized learning approaches like federated learning (FL) [4, 5] enable privacy-preserving collaborative model training. An FL system functions by collecting trained model parameters from edge devices (a.k.a., clients) rather than data. This mechanism safeguards the confidentiality of private data held by different parties. As a result, federated self-supervised learning (FedSSL) approaches have recently emerged as a potential strategy for leveraging raw data in distributed environments [6, 7, 8].
@@
Self-supervised learning (SSL), especially when involving state-of-the-art models like transformers [9, 10], can impose substantial computation demands. Unfortunately, clients in an FL environment often operate with limited computation and communication resources. Many clients may lack the necessary resources to participate in a conventional end-to-end FedSSL process. As a result, the data from these clients is excluded from the training process, potentially degrading the performance of the final model. Therefore, it is crucial to explore resource-efficient SSL approaches for FL systems.
A straightforward way to reduce computation demands on low-memory devices is to conduct the training with a smaller batch size. However, using a small batch size for SSL approaches can diminish the quality of learned representations and compromise performance [1, 7]. Moreover, simply reducing the batch size does not lower the FL communication cost since clients still need to exchange the full model with the server. On the other hand, using a large batch size would exclude low-memory devices from participating in the FL process.
@@
Table 1: FedMoCo (end-to-end) vs. FedMoCo-LW (layer-wise). We use a ViT-Tiny backbone with MoCo v3 as the SSL approach. The FL process involves 10 clients using a batch size of 1024 and the number of communication rounds set to 180. The unlabeled portion of the STL-10 dataset is used to setup the clients’ data, while the labeled training and testing portions are used for linear evaluation. We compare resource requirements for a single client.
@@
FedMoCo 15601 27.42 7352 54.99
FedMoCo-LW 3927 9.58 613 46.41
One promising solution is the layer-wise training approach, where a model is systematically trained one layer at a time, gradually increasing its depth. The term “layer” can refer to either an individual layer or a block comprising multiple layers. In layer-wise training, only a single layer is active at a time for training, significantly reducing computation requirements on FL clients. Moreover, only the active layer needs to be exchanged with the server, further reducing communication costs. Table 1 shows a significant reduction in a client’s resource requirements for layer-wise training compared to end-to-end training.
Although layer-wise training significantly reduces resource requirements, it may not be able to match the performance of end-to-end training, as shown in Table 1. In each stage of layer-wise training, a new layer is sequentially attached to the model, and all preceding layers are frozen, excluding them from further updates. Given a fixed number of FL rounds for each stage, it can be challenging to train the newly added layer to work seamlessly with the preceding layers, especially if the number of rounds is small. Although increasing the number of rounds could potentially alleviate this problem, it would impose extra communication and computation overhead on the clients, canceling the benefits of layer-wise training.
To address these challenges, we propose LW-FedSSL by introducing two simple, yet effective mechanisms: server-side calibration and representation alignment, designed to complement the layer-wise training process in an FL paradigm. Although clients in an FL environment may have limited resources, the central server typically possesses significant computation power. The server-side calibration mechanism leverages these resources to train the global model in an end-to-end manner, establishing layer cohesion. This mechanism solely operates on the central server and incurs no additional overhead on the clients. To preserve layer cohesion during local training, the representation alignment mechanism encourages alignment between local models and the server-side-trained global model.
Our contributions can be summarized as follows:
  • 1.
    We conduct an in-depth exploration of layer-wise training for SSL within the FL paradigm, considering various empirical aspects.
  • 2.
    We introduce LW-FedSSL, a novel layer-wise training framework designed to improve the resource efficiency of FL clients. By leveraging server-side calibration and representation alignment mechanisms, LW-FedSSL can compete with conventional end-to-end training counterpart while significantly lowering resource requirements.
  • 3.
    We explore Prog-FedSSL, a progressive training strategy. While this approach demands more resources than the layer-wise approach, it has the potential to enhance performance. Given its lower resource efficiency compared to LW-FedSSL, Prog-FedSSL is better suited for FL environments with more flexible resource constraints.
  • 4.
    We conduct a thorough evaluation of our proposed approaches through various FL settings and benchmarks.
2 Background
In this section, we briefly introduce fundamental concepts related to federated learning and self-supervised learning.
Figure 1: Four key steps in an FL communication round. (i) The server distributes a base global model $M$ to all participating clients. (ii) Each client $i\in[1,N]$ performs training on its local dataset $D^{i}$ to produce a local model $M^{i}$. (iii) Then, the local models are transmitted back to the server. (iv) The sever finally aggregates the received local models through weighted averaging to update the global model: $M\leftarrow\sum^{N}_{i=1}w^{i}M^{i}$ [5].
2.1 Federated Learning
A typical FL process involves a central coordinating server and a set of client devices, each of which stores private data [5]. The goal of FL is to train a global model $M$, by learning from the data residing on clients while ensuring their privacy. The FL objective can be expressed as
@@ 2.1 Federated Learning
where $N$ is the number of clients and $\mathcal{L}$ is the local loss. Here, $w^{i}=|D^{i}|/|D|$, which is the weight term of the $i$-th client, with $|D^{i}|$ denoting the number of samples in the local dataset $D^{i}$ and $|D|$ denoting the total number of samples in ${D=\bigcup_{i=1}^{N}D^{i}}$. As shown in Fig. 1, a single FL communication round consists of four main steps, which are repeated for a number of rounds $R$.
2.2 Self-Supervised Learning
Self-supervised learning (SSL) offers a means to utilize unlabeled data, which is often more readily available than labeled data for model training. SSL generates its own supervisory signals to learn valuable representations from the unlabeled data. We select MoCo v3 [11] as the SSL approach in our work. As shown in Fig 2, MoCo v3 features a Siamese network structure with an actively trained online branch, and a target branch, which is a moving-averaged copy of the online branch. The online branch consists of an encoder $F$, a projection head $H$, and a prediction head $P$. Meanwhile, the target branch consists of a momentum encoder $F_{k}$ and a momentum projection head $H_{k}$. Given an input sample $x$, it undergoes augmentation, creating two views, $x_{q}$ and $x_{k^{+}}$. These views are then fed into the online and target branches, respectively, to obtain a vector $q$ and its corresponding positive pair, $k^{+}$. Likewise, negative pairs, $k^{-}$, are derived from other samples within the same batch through the target branch. The online branch is trained using the InfoNCE loss [12],
@@ 2.2 Self-Supervised Learning
2.3 Federated Self-Supervised Learning
Although numerous studies on FL primarily focus on supervised learning, this means that device owners are tasked with labeling data, which is a tedious process. Consequently, client devices often lack labeled data due to limitations in data collection methods, labeling constraints, and insufficient annotation expertise and resources. In response, federated self-supervised learning (FedSSL) emerges as a solution to extend the applicability of FL to new domains where labeled data is scarce or difficult to acquire [13, 14, 15]. Many studies on FedSSL build on state-of-the-art SSL approaches like SimCLR [1], MoCo [2], and BYOL [16]. These studies adapt SSL techniques to address specific FL tasks [17, 18] and challenges at hand [19, 20, 21, 6]. In the context of FedSSL, the local dataset $D^{i}$ within a client $i$ is unlabeled, and thus, the local training process involves utilizing a form of SSL loss to train the local model $M^{i}$ on $D^{i}$. When using MoCo v3 as the SSL approach, this local model can comprise an encoder $F^{i}$ and its accompanying MLP heads, $H^{i}$ and $P^{i}$.
3 Related Work
Figure 2: Self-supervised learning with MoCo v3 [11]. We omit the negative samples for clarity.
@@ 3 Related Work
The study in [26] investigates layer-wise SSL in the speech domain. While it was intended for an on-device training scenario, the authors mainly examined it in a centralized setting. In [25], the authors propose an approach called federated layer-wise learning (FLL), based on SimCLR [1]. They introduce a depth dropout technique that randomly drops frozen layers during training to further reduce resource overhead. However, FLL is only examined with a large number of communication rounds in each layer-wise training stage (ranging from 4k to 12k), potentially placing significant strain on clients. Additionally, the exploration of FLL is limited in scope, missing comprehensive evaluation across diverse FL settings and benchmark datasets. In contrast to the above studies, we aim to offer a more in-depth exploration of layer-wise training in the context of FedSSL. Moreover, as we have previously demonstrated, layer-wise training alone may not achieve the same level of performance as end-to-end training. To bridge this gap, we introduce LW-FedSSL, a layer-wise training approach that can compete end-to-end training with significantly lower client resource requirements. One study [24] investigates the progressive training approach for FL known as ProgFed. Although similar to layer-wise training in gradually increasing the model depth, ProgFed differs by training all existing layers at each stage. However, ProgFed primarily focuses on federated supervised learning tasks that require labeled data for training. The nature of a progressive training approach within the FedSSL paradigm remains unexplored. To fill this gap, we introduce a progressive training strategy for FedSSL, referred to as Prog-FedSSL.
Figure 3: Layer-wise training process across different stages for the $i$-th client. Each stage $s\in[1,S]$ increases the depth of the encoder $F$ by adding a new layer $L_{s}$. During stage $s$, only the corresponding layer $L_{s}$ is actively updated, while prior layers (i.e., $L_{1}$ to $L_{s-1}$) are frozen.
4 Proposed Method
4.1 FedMoCo-LW
Firstly, we discuss the pure layer-wise training baseline, referred to as FedMoCo-LW. Fig. 3 illustrates the local training process of an FL client across different stages. For an encoder $F$ with a total of $S$ layers, the training process is divided into $S$ stages. Each stage $s\in[1,S]$ runs for a fixed number of FL communication rounds, $R_{s}$. The training process starts with an empty encoder $F$ without any layers, and each stage $s$ sequentially adds a new layer $L_{s}$ to $F$, gradually increasing its depth. During stage $s$, only the corresponding layer $L_{s}$ is updated, while all prior layers $L_{1}$ to $L_{s-1}$ within $F$ are frozen, serving only for inference. This strategy saves memory and computation, benefiting resource-constrained devices in FL environments. Moreover, in each communication round, clients are only required to send the active layer $L_{s}$ (and the MLP heads) to the server for aggregation. Likewise, the server only needs to broadcast the aggregated layer $L_{s}$ back to clients. This reduces both upload and download communication costs for clients. Although layer-wise training offers better resource efficiency than end-to-end training, there remains a notable performance gap in linear evaluation, as we have shown in Table 1.
4.2 LW-FedSSL
To address the performance gap between layer-wise and end-to-end training for FedSSL, we present LW-FedSSL by introducing two crucial mechanisms: server-side calibration, integrated at the server, and representation alignment, integrated into the local training process. Fig. 4 illustrates these two mechanisms within LW-FedSSL at stage $s$.
Server-side calibration. In layer-wise training, only the most recently added layer, $L_{s}$, is actively trained by clients at any given stage $s$. This may limit prior layers’ ability to adapt to the newly added layer, potentially degrading the model’s performance. Given a limited number of rounds $R_{s}$ available for each stage, solely training $L_{s}$ may prove insufficient to match the training received by the end-to-end strategy. On the other hand, increasing the number of rounds could introduce additional resource demands on clients. To tackle this, we can take advantage of resources available at the FL server. The server can contribute by training the current sub-model on a small auxiliary dataset, $D^{g}$, at the end of every communication round. Specifically, following the aggregation process, the server can train the aggregated encoder $F$ across all existing layers (i.e., $L_{1}$ to $L_{s}$) before broadcasting $F$ back to the clients. The auxiliary dataset $D^{g}$ may originate from a public dataset with a different data distribution than that of the clients’ data, as the server cannot directly access data within clients. Furthermore, $D^{g}$ may typically be unlabeled, requiring server-side training to also utilize an SSL approach. Training across all existing layers in the current sub-model plays a crucial role in calibrating the layers to function harmoniously. Also, our assumption regarding the auxiliary data at the server can be reasonably met since, in practice, the server may require a small dataset for various purposes, such as performance monitoring, pre-training [27, 7, 28], distillation [27], and other tasks in the FL context [29, 19, 30]. Particularly for FedSSL tasks, probing datasets may be necessary at the server to monitor and ensure smooth training progress.
@@ 4.2 LW-FedSSL
11:         $L_{s}^{i},H^{i},P^{i}\leftarrow\text{Local\_SSL}(i,F,H,P)$triangleright Algo. 2
@@
Representation alignment. FL clients often exhibit distinct local data distributions resulting from diverse device usage patterns and application settings. Such data heterogeneity can introduce bias to the local training process, negatively affecting the performance of the global model. To mitigate this, some studies [31, 32] encourage local training processes across different clients to converge toward parameters that are closely aligned with the global parameters. Often, this is achieved by aligning the representations of local models with those of the global model while simultaneously pushing them away from old representations (from old local models of previous FL rounds) [31, 33]. In contrast, our representation alignment mechanism does not require old local parameters. As a result, clients no longer need to store and perform inference on old parameters, reducing the overhead. Given a positive pair of views, $x_{q}$ and $x_{k^{+}}$, our approach uses local encoder $F^{i}$ and the global encoder $F$ to extract local representation $z^{i}=F^{i}(x_{q})$ and positive global representation $z^{+}=F(x_{k^{+}})$. Negative global representations $z^{-}$ are extracted from other samples. Similar to Equation 2, we can calculate the alignment loss as
The server-side calibration mechanism trains all existing layers in the global model to facilitate smooth collaboration between them. However, during local training, updates only to the active layer may disrupt this cohesion. To mitigate this, representation alignment encourages the updates of the local model to remain closely aligned with the parameters of the server-side trained global model. This helps alleviate complete overwriting and preserves the layer cohesion established by server-side calibration, while also allowing local models to learn new information from their local datasets. Additionally, representation alignment prevents extensive divergence between local models of different clients, enabling a more effective FL aggregation process.
In Section 5.3, we empirically show that server-side calibration and representation alignment are complementary for improving the effectiveness of layer-wise training. Integrating these mechanisms enables LW-FedSSL to compete with end-to-end training, all the while substantially reducing resource demands on FL clients. The details of LW-FedSSL are presented in Algorithms 1 and 2. Algorithm 1 outlines the server-side execution, which handles the gradual increase in the model depth and training stage, the server-side calibration mechanism, and the aggregation process. Each client executes Algorithm 2, performing local SSL training with representation alignment as regularization.
@@
Input: local dataset $D^{i}$, number of local epochs $E$, momentum $\mu$, temperature $\tau$, weight term $\alpha$
@@
1:Client executes: Local_SSL($i,F,H,P$):
2:Online branch: $F^{i}\leftarrow F$ and $H^{i}\leftarrow H$ and $P^{i}\leftarrow P$
3:Target branch: $F^{i}_{k}\leftarrow F$ and $H^{i}_{k}\leftarrow H$
4:for epoch $e=1,2,\dots,E$ do
5:     for each batch $x\in D^{i}$ do
6:         $x_{1}\leftarrow\text{{Augment}}(x)$ and $x_{2}\leftarrow\text{{Augment}}(x)$
7:         $z_{1}\leftarrow F(x_{1})$ and $z_{2}\leftarrow F(x_{2})$
8:         $z_{1}^{i}\leftarrow F^{i}(x_{1})$ and $z_{2}^{i}\leftarrow F^{i}(x_{2})$
9:         $q_{1}^{i}\leftarrow P^{i}(H^{i}(z_{1}^{i}))$ and $q_{2}^{i}\leftarrow P^{i}(H^{i}(z_{2}^{i}))$
10:         $k_{1}^{i}\leftarrow H^{i}_{k}(F^{i}_{k}(x_{1}))$ and $k_{2}^{i}\leftarrow H^{i}_{k}(F^{i}_{k}(x_{2}))$
11:         $\ell_{\text{con}}\leftarrow\ell(q^{i}_{1},k^{i}_{2},\tau)+\ell(q^{i}_{2},k^{i}_{1},\tau)$triangleright Eq. (2)
12:         $\ell_{\text{align}}\leftarrow\ell(z^{i}_{1},z_{2},\tau)+\ell(z^{i}_{2},z_{1},\tau)$triangleright Eq. (3)
13:         $\mathcal{L}\leftarrow\ell_{\text{con}}+\alpha\ell_{\text{align}}$
14:         $F^{i},H^{i},P^{i}\leftarrow\text{Update}(\nabla\mathcal{L})$
15:         $F^{i}_{k},H^{i}_{k}\leftarrow\text{Momentum\_Update}(F^{i},H^{i},\mu)$
16:$L_{s}^{i}\leftarrow\text{Get active layer from }F^{i}$
17:return $L_{s}^{i},H^{i},P^{i}$
Table 2: Given an encoder $F$ with a total of $S$ layers, we compare the characteristics of different training approaches at stage $s\in[1,S]$.
@@
FedMoCo-LW layer-wise $S$ $L_{s}$ $L_{1}$ to $L_{s-1}$ $L_{s}$ $L_{s}$ $L_{s}$